There are two types of canned coffee with the same volume, milk coffee and black coffee. The milk and water in milk coffee is 1:4 with a volume of 375ml. The milk in black coffee is 50% less than the milk in milk coffee. Find the volume of milk to coffee in black coffee.
Answer by stanbon(75887)

There are two types of canned coffee with the same volume, milk coffee and black coffee.
The milk and water in milk coffee is 1:4 with a volume of 375ml.
Let "m" be the amount of milk; Then "4m" is the amount of water.
Equation:: m + 4m = 375ml; 5m= 375ml: m = 75ml (amt. of milk in milk coffee)
----------------------
The milk in black coffee is 50% less than the milk in milk coffee.
Find the volume of milk to coffee in black coffee.
Equation:
Ans: amt. of milk in black coffee = 0.5*375 = 187.5 ml
-----------
Cheers,
Stan H.
----------